# Superconductivity from spin-canting fluctuations in rhombohedral graphene

### v3

@@ I Introduction
Our scenario also accommodates more refined experimental observations enumerated earlier. First, recall that spin-orbit-enabled superconductivity entails minority Fermi pockets; moreover, in trilayers superconductivity abruptly turns off when they are vacated through a Lifshitz transition [44]. These observations hint that pairing requires the presence of minority carriers, in harmony with spin-flipping scattering that acts between the majority and minority Fermi surfaces. Second, Ref. 42 observed a tunable and gradual increase of the maximal $T_{c}$ with induced Ising spin-orbit-coupling $\lambda_{I}$, coinciding with a reduction in the phase space occupied by superconductivity. The concurrence of these two trends is surprising because in a conventional scenario, where the pairing glue is non-magnetic (e.g., phonon-mediated), spin-orbit coupling is expected to either uniformly help or hurt superconductivity through its effect on the Fermi surface. Our scenario, however, naturally recovers these seemingly at odds trends: Increasing $\lambda_{I}$ decreases the canting angle $\theta_{0}$ [see Eq. (4)]—in turn enhancing the magnon-mediated pairing interaction [see Eq. (38)] and thereby boosting $T_{c}$. At the same time, spin-orbit interaction competes against Hund’s coupling, such that increasing $\lambda_{I}$ erodes the phase space occupied by the spin-canted phase, whose Goldstone modes provide the pairing glue in our scenario. Increasing $\lambda_{I}$ thus also narrows the superconducting phase. Recent Hartree-Fock simulations support the predicted change in $\theta_{0}$ with $\lambda_{I}$ [54] accounting for the rise in $T_{c}$, while also finding that the phase space for spin-canting order evolves with $\lambda_{I}$ similarly to the phase space for superconductivity in experiment (see Fig. 7 in Ref. 53 and Fig. 4 in Ref. 54). To the best of our knowledge, no other pairing mechanism relate superconductivity to spin canting in these systems.
@@ II.1 Landau-theory description
In the first term—which is invariant under independent SU(2) spin rotations in the two valleys—$\kappa_{s}$ encodes a combination of intravalley exchange interactions and kinetic energy. Near the onset of a Stoner transition driven primarily by these ingredients, $\kappa_{s}$ turns negative and promotes the development of non-zero order parameters ${\boldsymbol{n}}_{\xi}$. Intervalley exchange processes are modeled by the second term, which breaks the symmetry of the $\kappa_{s}$ term down to global SU(2) spin rotations. We assume $J>0$, corresponding to ferromagnetic Hund’s coupling that favors aligning spins for the two valleys. In the third term, $\tau_{K,K^{\prime}}=\pm 1$ is a valley-dependent sign and $\lambda_{I}$ denotes Ising spin-orbit coupling that favors spins pointing out-of-plane, in opposite directions in the two valleys, denoted as ±$\pm\hat{\bf z}$. Ising spin-orbit coupling further reduces the symmetry down to global U(1) spin rotations around the $z$ axis. For concreteness, we assume $\lambda_{I}\geq 0$ throughout. Finally, the ellipsis denotes symmetry-allowed that are higher-order in ${\boldsymbol{n}}_{\xi}$ (e.g., cubic and quartic terms). Such terms fix the order-parameter magnitudes in the ordered phase and also determine whether the Stoner transition is continuous or first-order [57]. Importantly, our analysis applies regardless of the order of the phase transition.
@@ II.4 Magnon fluctuations
where $\kappa=[4(V+J)\bar{m}^{2}]^{-1}$, $D$ is a non-universal constant, and $\gamma=2\kappa\bar{m}$. Note that we ignored the self-energy in the electron Green’s function, which can shift the dispersion (real part) and blur the Fermi surface (imaginary part). This omission is justified because the self-energy is not expected to alter the qualitative features of the magnons—e.g., their degeneracy, and whether their dispersion is linear or quadratic—that are dictated by general symmetry considerations [59].
With this preparation, we proceed to work out the magnon dispersion. Hund’s coupling $J$ embedded in the first line of Eq. (II.4) hybridizes fluctuations in the two valleys. Consequently, it is convenient to pass to a basis of symmetric and antisymmetric valley fluctuations via
@@ III.1 Canting angles $\theta_{0}<\pi/2$
The first term on the right-hand side arises from the $s-$ Goldstone mode, which remains gapless throughout the ordered phase, while the second term arises from the low-lying $a-$ mode that becomes gapless at criticality. As a reminder, the $a-$ magnon gap is smaller than $\epsilon_{F}$ by an $\mathcal{O}(J/V)$ factor; see text below Eq. (30). As a result, pair scattering mediated by this mode features significant retardation, and therefore will be roughly as relevant for pairing as the Goldstone mode. At the critical point where $\Omega_{a}\rightarrow 0$, the restored U(1) symmetry enforces a degeneracy of the two gapless modes, which manifests as $c_{s}=c_{a}$ and $z_{s}=z_{a}$.
@@ IV Pairing from magnons
$\Delta_{+}(k)=-\sum_{k^{\prime}}g^{-+,-+}_{K,K^{\prime}}(k-k^{\prime})\frac{\Delta_{-}(k^{\prime})}{\omega_{n}^{\prime 2}+\epsilon_{{\boldsymbol{k}}^{\prime},-}^{2}},$ (40)
$\Delta_{-}(k)=-\sum_{k^{\prime}}g^{+-,+-}_{K,K^{\prime}}(k-k^{\prime})\frac{\Delta_{+}(k^{\prime})}{\omega_{n}^{\prime 2}+\epsilon_{{\boldsymbol{k}}^{\prime},+}^{2}},$ (41)
@@
To proceed analytically, we apply the trick conventionally used in solving the BCS problem: replacing the full pairing interaction with a frequency-independent version that exhibits a hard cut-off at the characteristic magnon energy $\Omega_{0}$ (analogous to the Debye energy in conventional phonon-mediated pairing). For simplicity, we focus on the small canting angle regime so that, up to first order in $J/V$, the two magnon velocities are similar, $c_{s}\sim c_{a}\sim c=\sqrt{\frac{JD}{V\kappa}}$, and $z_{s}\sim z_{a}\sim z=\frac{V}{2J}$. Here the characteristic magnon energy is estimated as $\Omega_{0}=ck_{0}$, where $k_{0}$ is a characteristic scale for the momentum transfer ${\boldsymbol{k}}-{\boldsymbol{k}}^{\prime}$ in the spin-flipping scattering process. This scale is upper-bounded by $2k_{F}^{+}$, with $k_{F}^{+}$ the Fermi momentum for the majority Fermi surface. However, in rhombohedral graphene the momentum-space separation between majority and minority pockets can be several times smaller than $k_{F}^{+}$, due in particular to trigonal warping (e.g., Ref. 53).
@@
$\frac{1}{(\omega_{n}-\omega^{\prime}_{n})^{2}+c^{2}(k-k^{\prime})^{2}}$ $\rightarrow\frac{\Theta(\Omega_{0}-|\omega_{n}^{\prime}|)\Theta(\Omega_{0}-|\omega_{n}|)}{\Omega_{0}^{2}},$
$\Delta_{\pm}(\omega_{n},k)$ $\rightarrow\Delta_{\pm}^{(0)}\Theta(\Omega_{0}-|\omega_{n}|),$ (42)
@@
$\Delta_{-}^{(0)}=-g_{\rm eff}\nu_{+}\Delta_{+}^{(0)}\log\frac{\Omega_{0}}{T_{c}},$ (44)
with an effective pairing strength
$g_{\rm eff}\approx\frac{\cos^{2}\theta_{0}}{\frac{V}{2J}\kappa\Omega_{0}^{2}}=\frac{2\cos^{2}\theta_{0}}{Dk_{0}^{2}}.$ (45)
@@
This result enables a crude $T_{c}$ estimate. The dimensionless coupling constant $g_{\rm eff}\nu_{\rm eff}$ is independent of the ratio $J/V$ and only involves band structure quantities. Therefore, it is expected to be of $\mathcal{O}(1)$ from dimensional analysis, but will depend on the fermiology and density of states of a specific system. The magnon energy $\Omega_{0}$ is proportional to a small dimensionless ratio $\sqrt{J/V}$ times the Fermi energy (again assuming that $\bar{m}\sim\epsilon_{F})$. In realistic systems, we expect $V/J\sim O(10)$. Therefore, the superconducting $T_{c}$ predicted in our scenario can reach a substantial fraction of $\epsilon_{F}$. Given that we neglected the competing effect of repulsive Coulomb interactions, this estimate should be understood as an upper bound on $T_{c}$. Indeed, in lightly doped rhombohedral bilayers and trilayers (where $\epsilon_{F}\sim 1-10$ meV), the reported critical temperatures [40, 42, 41, 44, 43, 45] for the spin-orbit enabled superconductors are only in the range $k_{B}T_{c}\sim 10-50\mu$eV.
@@ V Discussion
Our pairing scenario also has ramifications for depairing by in-plane magnetic fields $B_{\parallel}$. Conventional in-plane-field depairing mechanisms include orbital effects and warping of the Fermi surfaces due to an interplay with Rashba spin-orbit coupling [66] (which we have neglected so far but discuss further below). Since the pairing glue in our scenario is based on magnons, one might suspect a non-trivial magnetic-field dependence on $T_{c}$. Indeed, $B_{\parallel}$ significantly affects the magnon spectrum near ${\boldsymbol{q}}=0$ as it explicitly breaks the U(1) spin-rotation symmetry that the spin-canting phase breaks spontaneously—thereby generating a gap $\Delta_{\rm mag}$ for the Goldstone magnon mode. However, since our pairing mechanism utilizes magnons near $|{{\boldsymbol{q}}}|\sim k_{0}$ rather than ${{\boldsymbol{q}}}=0$, $B_{\parallel}$ leads to a weak suppression of the pairing glue strength, until the induced magnon gap $\Delta_{\rm mag}$ becomes comparable to the typical magnon energy $\Omega_{0}$ at $|{{\boldsymbol{q}}}|\sim k_{0}$. While our mechanism implies a greater depairing effect of in-plane magnetic fields compared to, say, a phonon-mediated scenario, its effect in a given platform will depend on whether orbital depairing, Rashba or magnon effects are predominant.
@@
We conclude with some speculation on extensions of our theory. Analyzing the favorability of spin-canting order in systems composed of more than three graphene layers, e.g., from a Hartree-Fock angle, would provide valuable insight into the wider applicability of our proposed pairing mechanism. It is also interesting to explore pairing mediated by Goldstone modes associated with other broken-symmetry orders [67], most notably intervalley coherence that also naturally emerges in graphene multilayers [12, 21, 50, 68, 53, 69, 70, 54, 71, 72, 73, 74]. A particularly interesting possibility—which combines the two different types of continous symmetry breaking relevant for graphene systems—consists of intervalley coherent orders that also exhibit spin canting [54]. Such composite order can lead to generalized quarter-metal states that comprise a single, non-degenerate majority Fermi surface; similarly to the ideas developed in this work, our magnon-mediated pairing mechanism could potentially mediate Cooper pairing in this setting. Interestingly, one of the new superconducting regions in proximitized Bernal bilayer graphene [40] exhibits fermiology consistent with this description.

### v4

@@ I Introduction
Our scenario also accommodates more refined experimental observations enumerated earlier. First, recall that spin-orbit-enabled superconductivity entails minority Fermi pockets; moreover, in trilayers superconductivity abruptly turns off when they are vacated through a Lifshitz transition [44]. These observations hint that pairing requires the presence of minority carriers, in harmony with spin-flipping scattering that acts between the majority and minority Fermi surfaces. Second, Ref. 42 observed a tunable and gradual increase of the maximal $T_{c}$ with induced Ising spin-orbit-coupling $\lambda_{I}$, coinciding with a reduced (yet still extended) phase space occupied by superconductivity. The concurrence of these two trends is surprising because in a conventional scenario, where the pairing glue is non-magnetic (e.g., phonon-mediated), spin-orbit coupling is expected to either uniformly help or hurt superconductivity through its effect on the Fermi surface. Our scenario, however, naturally recovers these seemingly at odds trends: Increasing $\lambda_{I}$ decreases the canting angle $\theta_{0}$ [see Eq. (4)]—in turn enhancing the magnon-mediated pairing interaction [see Eq. (38)] and thereby boosting $T_{c}$. At the same time, spin-orbit interaction competes against Hund’s coupling, such that increasing $\lambda_{I}$ erodes the phase space occupied by the spin-canted phase, whose Goldstone modes provide the pairing glue in our scenario. Increasing $\lambda_{I}$ thus also narrows the superconducting phase. Recent Hartree-Fock simulations support the predicted change in $\theta_{0}$ with $\lambda_{I}$ [54] accounting for the rise in $T_{c}$, while also finding that the phase space for spin-canting order evolves with $\lambda_{I}$ similarly to the phase space for superconductivity in experiment (see Fig. 7 in Ref. 53 and Fig. 4 in Ref. 54). To the best of our knowledge, no other pairing mechanism relates superconductivity to spin canting in these systems.
@@ II.1 Landau-theory description
In the first term—which is invariant under independent SU(2) spin rotations in the two valleys—$\kappa_{s}$ encodes a combination of intravalley exchange interactions and kinetic energy. Near the onset of a Stoner transition driven primarily by these ingredients, $\kappa_{s}$ turns negative and promotes the development of non-zero order parameters ${\boldsymbol{n}}_{\xi}$. Intervalley exchange processes are modeled by the second term, which breaks the symmetry of the $\kappa_{s}$ term down to global SU(2) spin rotations. We assume $J>0$, corresponding to ferromagnetic Hund’s coupling that favors aligning spins for the two valleys. In the third term, $\tau_{K,K^{\prime}}=\pm 1$ is a valley-dependent sign and $\lambda_{I}$ denotes Ising spin-orbit coupling that favors spins pointing out-of-plane, in opposite directions in the two valleys, denoted as ±$\pm\hat{\bf z}$. Ising spin-orbit coupling further reduces the symmetry down to global U(1) spin rotations around the $z$ axis. For concreteness, we assume $\lambda_{I}\geq 0$ throughout. Finally, the ellipsis denotes symmetry-allowed terms that are higher-order in ${\boldsymbol{n}}_{\xi}$ (e.g., cubic and quartic terms). Such terms fix the order-parameter magnitudes in the ordered phase and also determine whether the Stoner transition is continuous or first-order [57]. Importantly, our analysis applies regardless of the order of the phase transition.
@@ II.4 Magnon fluctuations
where $\kappa=[4(V+J)\bar{m}^{2}]^{-1}$, $D$ is a non-universal constant, and $\gamma=2\kappa\bar{m}$. Note that we ignored the self-energy in the electron Green’s function, which can shift the dispersion (real part) and blur the Fermi surface (imaginary part). This omission is justified because the self-energy is not expected to alter the qualitative features of the magnons—e.g., their degeneracy, and whether their dispersion is linear or quadratic—that are dictated by general symmetry considerations [59]. With this preparation, we proceed to work out the magnon dispersion. Hund’s coupling $J$ embedded in the first line of Eq. (II.4) hybridizes fluctuations in the two valleys. Consequently, it is convenient to pass to a basis of symmetric and antisymmetric valley fluctuations via
@@ III.1 Canting angles $\theta_{0}<\pi/2$
The first term on the right-hand side arises from the $s-$ Goldstone mode, which remains gapless throughout the ordered phase, while the second term arises from the low-lying $a-$ mode that becomes gapless at criticality. As a reminder, the $a-$ magnon gap is smaller than $\epsilon_{F}$ by an $O(J/V)$ factor; see text below Eq. (30). As a result, pair scattering mediated by this mode features significant retardation, and therefore will be roughly as relevant for pairing as the Goldstone mode. At the critical point where $\Omega_{a}\rightarrow 0$, the restored U(1) symmetry enforces a degeneracy of the two gapless modes, which manifests as $c_{s}=c_{a}$ and $z_{s}=z_{a}$.
@@ IV Pairing from magnons
$\Delta_{+}(k)=-\sum_{k^{\prime}}g^{-+,-+}_{K,K^{\prime}}(k-k^{\prime})\frac{\Delta_{-}(k^{\prime})}{\omega^{\prime 2}+\epsilon_{{\boldsymbol{k}}^{\prime},-}^{2}},$ (40)
$\Delta_{-}(k)=-\sum_{k^{\prime}}g^{+-,+-}_{K,K^{\prime}}(k-k^{\prime})\frac{\Delta_{+}(k^{\prime})}{\omega^{\prime 2}+\epsilon_{{\boldsymbol{k}}^{\prime},+}^{2}},$ (41)
@@
To proceed analytically, we apply the trick conventionally used in solving the BCS problem: replacing the full pairing interaction with a frequency-independent version that exhibits a hard cut-off at the characteristic magnon energy $\Omega_{0}$ (analogous to the Debye energy in conventional phonon-mediated pairing). We first focus on the critical point, $\theta_{0}=0$, where pairing is the strongest. Here the two magnon velocities are identical by symmetry, $c_{s}=c_{a}=c$, with $c\approx\sqrt{\frac{2JD}{V\kappa}}$ in the limit $V\gg J$. The characteristic magnon energy is estimated as $\Omega_{0}=ck_{0}$, where $k_{0}$ is a characteristic scale for the momentum transfer ${\boldsymbol{k}}-{\boldsymbol{k}}^{\prime}$ in the spin-flipping scattering process. This scale is upper-bounded by $2k_{F}^{+}$, with $k_{F}^{+}$ the Fermi momentum for the majority Fermi surface. However, in rhombohedral graphene the momentum-space separation between majority and minority pockets can be several times smaller than $k_{F}^{+}$, due in particular to trigonal warping (e.g., see Ref. 53).
@@
$\frac{1}{(\omega-\omega^{\prime})^{2}+c^{2}({\boldsymbol{k}}-{\boldsymbol{k}}^{\prime})^{2}}$ $\rightarrow\frac{\Theta(\Omega_{0}-|\omega^{\prime}|)\Theta(\Omega_{0}-|\omega|)}{\Omega_{0}^{2}},$
$\Delta_{\pm}(\omega,{\boldsymbol{k}})$ $\rightarrow\Delta_{\pm}^{(0)}\Theta(\Omega_{0}-|\omega|),$ (42)
@@
$\Delta_{-}^{(0)}=-g_{\rm eff}\nu_{+}\Delta_{+}^{(0)}\log\frac{\Omega_{0}}{T_{c}}.$ (44)
Here, $g_{\rm eff}$ is the pair scattering amplitude evaluated at the smallest momentum connecting the two Fermi surfaces: $g_{\rm eff}=g_{K,K^{\prime}}^{-+,-+}(k_{0})$ with $k_{0}=k_{F+}-k_{F-}$.
At the canting critical point $\theta_{0}=0$, the $a-$ mode is gapless and degenerate with the $s-$ mode. Both modes contribute constructively and the pairing interaction is maximized, with
$g_{\rm eff}\approx\frac{4J}{V\kappa\Omega_{0}^{2}}\approx\frac{2}{Dk_{0}^{2}}.$ (45)
Equation (45) holds when the magnons used in the pair-scattering process disperse linearly. This assumption holds below a momentum cutoff $q_{\rm max}\sim O(\sqrt{J/V})k_{F}$. (At larger $|{{\boldsymbol{q}}}|$, magnons disperse quadratically; Appendix D provides more general expressions for scattering amplitudes valid also in that regime.) Our assumption that pair scattering transfers a momentum $|{{\boldsymbol{q}}}|\lesssim q_{\rm max}$ is reasonable because (1) $V/J$ is expected to be $\sim O(10)$ in rhombohedral graphene devices (see Appendix C), so that $\sqrt{V/J}\sim 3$ is not too large; and (2) due to trigonal warping, pair scattering between minority and majority pockets can occur with momenta a few times smaller than $k_{F}$.
@@
This result enables a crude $T_{c}$ estimate near the critical point. The magnon energy $\Omega_{0}$ is proportional to a small dimensionless ratio $\sqrt{J/V}$ times the Fermi energy (again assuming that $\bar{m}\sim\epsilon_{F})$. Appendix C provides an experimentally motivated discussion that estimates $\Omega_{0}\sim 1$ meV. The dimensionless coupling constant $g_{\rm eff}\nu_{\rm eff}$ is independent of the ratio $J/V$ and only involves band structure quantities—it is expected to be of $O(1)$ from dimensional analysis, but will depend on the fermiology and density of states of a specific system. The superconducting $T_{c}$ predicted in our scenario can therefore reach a substantial fraction of $\Omega_{0}$. Given that we neglected the competing effect of repulsive Coulomb interactions, this estimate should be understood as an upper bound on $T_{c}$. Indeed, in lightly doped rhombohedral bilayers and trilayers the reported critical temperatures [40, 42, 41, 44, 43, 45] for the spin-orbit enabled superconductors are only in the range $k_{B}T_{c}\sim 10-50~{}\mu$eV.
Moving away from the critical point, pairing is suppressed by the $\cos^{2}\theta_{0}$ numerator in Eq. (38) and the opening of a gap in the anti-symmetric mode. For a generic $\theta_{0}$, we have
$g_{\rm eff}\approx\frac{2J\cos^{2}\theta_{0}}{V\kappa}\left(\frac{1}{c_{s}^{2}k_{0}^{2}}+\frac{1}{\Omega_{a-}^{2}+c_{a}^{2}k_{0}^{2}}\right).$ (47)
The first and second terms respectively correspond to the contributions of the $s-$ and $a-$ modes. Here the cutoff frequency scale $\Omega_{0}$ should be determined separately for the two magnon modes: $\Omega_{0}\sim c_{s}k_{0}$ for the $s-$ mode, and $\Omega_{0}\sim\sqrt{\Omega_{a-}^{2}+c_{a}^{2}k_{0}^{2}}$ for the $a-$ mode. However, when $\theta$ is not too close to $\pi/2$ these two frequencies are comparable and a few times below the Fermi energy [^3] Our numerics in Appendix C check this claim and show that the magnon energy is of order $\sim 1$meV at $q_{\rm{max}}$, indeed a few times smaller than $E_{F}$., such that the pairing interaction exhibits some retardation. At $\theta_{0}=\pi/2$, the pairing interaction vanishes due to the symmetry restrictions higlighted in Sec. III.2.
@@ V Discussion
Our pairing scenario also has ramifications for depairing by in-plane magnetic fields $B_{\parallel}$. Conventional in-plane-field depairing mechanisms include orbital effects and warping of the Fermi surfaces due to an interplay with Rashba spin-orbit coupling [66] (which we have neglected so far but discuss further below). Since the pairing glue in our scenario is based on magnons, one might suspect a non-trivial magnetic-field dependence of $T_{c}$. Indeed, $B_{\parallel}$ significantly affects the magnon spectrum near ${\boldsymbol{q}}=0$ as it explicitly breaks the U(1) spin-rotation symmetry that the spin-canting phase breaks spontaneously—thereby generating a gap $\Delta_{\rm mag}\sim\sqrt{|B_{\parallel}|}$ for the Goldstone magnon mode (see Appendix E for a derivation). Interestingly, this $\sqrt{B_{\parallel}}$ magnon gap results in an unconventional behavior of the critical temperature reduction, which scales as $|B_{\parallel}|$ at weak fields instead of the traditional Ginzburg-Landau expectation $T_{c}-T_{c}^{(0)}\sim B_{\parallel}^{2}$. At the canting transition, the gap scaling is modified to $\Delta_{\rm mag}\sim|B_{\parallel}|^{1/3}$—resulting in an unusual power-law for depairing, with $T_{c}-T_{c}^{(0)}\sim B_{\parallel}^{2/3}$. Magnon-induced depairing is therefore most effective at the canting transition (where pairing is strongest) because of the enhanced susceptibility towards opening a magnon gap at weak fields. These features constitute unique experimental predictions of our pairing scenario. We note that while our mechanism implies a greater depairing effect of in-plane magnetic fields compared to, say, a phonon-mediated scenario, its observability in a given material platform will depend on whether orbital depairing, Rashba or magnon effects are predominant. Nevertheless, the dependence of the magnon gap on magnetic field and spin-orbit coupling could be imaged experimentally using, e.g., microwave resonator techniques [67, 68, 69].
@@
The soft magnons mediating Cooper pairing in this work are subject to Landau damping due to magnon decay into electron-hole pairs near the Fermi surface. A legitimate concern is whether such Landau damping—ignored so far in our analysis—will wash out the magnon-mediated pairing interaction. In Appendix F we find that Landau damping in our model modifies the spin susceptibility $\chi_{1}$ by a term linear in $|\nu|$, thereby suppressing the pairing interaction. However, since its impact is $O(|\nu|)$ and vanishes at $\nu=0$, there always exists a frequency scale—which we estimate through dimensional analysis to be of $O(\frac{J}{V}E_{F})$—below which the impact of damping is subleading. Given that the pairing mechanism mainly uses magnons with frequency below $O(\frac{J}{V}E_{F})$, we conclude that the impact of damping on the pairing interaction will be limited and that our analysis of pairing remains valid.
Throughout our work we also neglected the (material-dependent) form factors arising from projecting the pairing interaction into the relevant electronic bands. While these form factors carry some momentum dependence, our (projected) intervalley pairing interaction will remain of the same sign along the Fermi surface due to orbital time-reversal symmetry. We therefore do not expect the pairing symmetry to be altered: generally, introducing nodes in the gap function costs condensation energy, and is only preferred if the pairing interaction has a sign alternation along the Fermi surface.
We emphasize that our scenario aims to explain the appearance of spin-orbit-enabled superconductivity in TMD-proximitized devices, where the normal state is valley balanced. We therefore do not claim any relation to the intriguing superconducting phase found in quarter-metal regions of the penta-, tetra- and hexa-layer graphene phase diagrams [70, 71], where spin and valley are assumed to be both polarized. Neither do we claim a connection to the superconductivity found in pristine bilayer and trilayer graphene (without a TMD layer). Our perspective, rooted in recent measurements [42, 45, 44], is that superconductivity in graphene with and without a TMD substrate likely arises from different mechanisms. Indeed, superconductivity in TMD-proximitized devices emerges in regions of the phase diagram where it is absent in “bare” devices—whereas the superconducting phases present in the bare devices seem to be suppressed (or completely absent) with a TMD substrate.
We conclude with some speculation on extensions of our theory. Analyzing the favorability of spin-canting order in systems composed of more than three graphene layers, e.g., from a Hartree-Fock angle, would provide valuable insight into the wider applicability of our proposed pairing mechanism. It is also interesting to explore pairing mediated by Goldstone modes associated with other broken-symmetry orders [72], most notably intervalley coherence that also naturally emerges in graphene multilayers [12, 21, 50, 73, 53, 74, 75, 54, 76, 77, 78, 79, 80]. A particularly interesting possibility—which combines the two different types of continuous symmetry breaking relevant for graphene systems—consists of intervalley coherent orders that also exhibit spin canting [54]. Such composite order can lead to generalized quarter-metal states that comprise a single, non-degenerate majority Fermi surface; similarly to the ideas developed in this work, our magnon-mediated pairing mechanism could potentially mediate Cooper pairing in this setting. Interestingly, one of the new superconducting regions in proximitized Bernal bilayer graphene [40] exhibits fermiology consistent with this description.